the length of a rectangle is 7 more than the width. the area is 330 square centimeters. find the length and the width of the rectangle.
Answer by mananth(16946)

let width be x
length =x+7
..
Area = L*W
x(x+7)=330
x^2+7x=330
x^2+7x-330=0
x^2+22x-15x-330=0
x(x+22)-15(x+22)=0
(x+22)(x-15)=0
x= 15 is the width.
length = x+7 = 22